OpenStudy (anonymous):

Could somebody please check if I got this right Integral of cos(x)*e^sin(x) u=sin(x) du=cos(x) dx Integral cos(x)*cos(x)*e^u dx 1/cos(x) integral cos(x)*e^u du e^sin(x) +c

OpenStudy (anonymous):

the answer is ok, the process is a little bit not correct, hihihi

OpenStudy (anonymous):

Ok could you please show me how

OpenStudy (anonymous):

It probably should of been integral cos(x)*e^u dx

OpenStudy (anonymous):

u = sinx --> du = cosx dx, \[\int cosx e^{sinx}dx = \int e^{sinx}cosx dx\] replace the substitute in, you have \[\int e^u du\] that gives you the answer , not as you did

OpenStudy (anonymous):

e^u du = e^sin(x) + c

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Your process is not correct, re -read, please. YOu must understand what is going on to get the correct answer.